The denominator is 5 times the numerator. The sum of the numerator and the denominator is 36. Find the fraction.
Answer by ptaylor(1894)

Let x=numerator
Then 5x=denominator
We are told that:
x+5x=36 combine like terms
6x=36 divide both sides by 6
x=6--------------numerator
5x=5*6=30 denominator
x/5x=6/30-------------------------the fraction
CK
6+30=36
36=36
Hope this helps---ptaylor